# Solving Robotics Problems in Zero-Shot with Vision-Language Models

### v3

@@ 6 Further Discussions
6.1 Comparison with Methods that Train
In recent years, the machine learning community has often seen new LLMs exceed the performance of previous generation fine-tuned models in zero-shot settings, despite the latter’s advantage of task-specific tuning. To explore this trend in the context of visual LLMs and robotics, we compare Wonderful Team with several methods that were at least partially fine-tuned on robotics tasks.
@@ Acknowledgments
  • Appendix B: Experimental Details B
  • Appendix C: Ablations: Are All Parts of Wonderful Team Necessary? C
    • Understanding What Each Part of Wonderful Team Does C.1
  • Appendix D: Ablation Studies: VLLMs’ Spatial Reasoning Limitations and Potentials D
    • Evaluating VLLM’s Spatial Understanding D.1
    • Evaluating VLLMs’ Error Recognition and Correction D.2
  • Appendix E: Comparison with Other Methods E
    • Natural Language as Policies E.1
    • PIVOT E.2
    • Language Models as Zero-Shot Trajectory Generators E.3
@@ Appendix A Things are Moving Extremely Fast
In the Appendix D, we examine the impact of this rapid progress on the grounding team in particular, and show that older VLLMs often struggled to draw bounding boxes with any regularity, suggesting they lacked the fidelity needed for fine-grained robotic control.
Appendix B Experimental Details
B.1 Evaluation Protocol
@@
B.2 Multimodal Reasoning - Simulated VIMABench
@@ B.2.1 Task Details
B.2.2 Full Experimental Results
@@
Task Num VIMA 200M Instruct2Act NLaP (w/o CoT) NLaP TG Ours
1: Visual Manipulation 99 91 93 100 60 100
2: Scene Understanding 100 81 60 67 40 100
3: Rotate 100 98 93 93 80 100
4: Rearrange 97 79 52 73 - 80
5: Rearrange then Restore 54.5 72 25 73 - 70
6: Novel Adjective 100 82 13 43 10 70
7: Novel Noun 99 88 8 80 0 100
*8: Novel Adjective and Noun - - - - - 60
*9: Twist 17.5 - - - - 50
10: Follow Motion - 35 0 12 - 10
11: Follow Order 90.5 72 0 0 - 0
12: Without Exceeding 93 68 17 47 10 90
*13: Without Touching - 0 0 3 - 40
14: Same Texture - 80 3 71 - 100
15: Same Shape 97.5 78 10 80 0 100
16: Manipulate Old Neighbor 46 64 8 20 0 90
17: Pick in Order then Restore 43.5 85 10 30 0 90
Tasks marked with a star were excluded from the main paper’s results for the following reasons: 1. Nature of Tasks: Categories Visual Goal Reaching (Task 4 and 5) and One-shot Video Imitation (Task 10 and 11) were excluded because these tasks are not the best indicators of VLLM’s capabilities without additional prompting.
@@
For example, as shown in Figure 13, Task 11 in the One-shot Video Imitation category requires examining several consecutive frames as ’goal scenes’. Without further task-specific prompting or training, it is very challenging to infer the required actions between frames since there isn’t a single correct answer. For instance, transitioning from Frame 1 to Frame 2 in this example could be achieved by moving the yellow O onto the red O, or by first removing the red O and then moving the yellow O to the same position. By nature, these tasks require additional tools or workflows, which complicate zero-shot evaluation. Additional prompting on tasks like this to help the VLLMs better understand the relationship between frames will probably be helpful. However, this is not the focus of our research, so we used the same prompt for these evaluations in Table 2.
2. Missing Baseline Results: Tasks 8, 9, 13, and 14 were excluded due to the lack of available baseline results for comparison.
A complete list of tasks with video illustrations can be found here.
B.3 Implicit Goal Inference - Real Robots
B.3.1 Task Details
@@ B.3.2 Robot Setup
B.3.3 Results
Our real robot experiments demonstrated that our framework successfully completed all three tasks 100% of the time. Note that we did not modify any of the prompt or pipeline moving from simulated VIMABench environment to the tasks on the real robot. It was surprising to us how robust the reasoning and planning capabilities of Wonderful Team are. This section provides qualitative results from these experiments, illustrated in Figures 15, 16, and 17. These figures highlight specific aspects of the tasks, illustrating the effectiveness of our framework. It is important to note that these results only reflect the work of the planning team. The role of the grounding team, locating objects and determining their positions, is crucial for the successful execution of these plans.
@@ B.3.3 Results
In the fruit placement task (Figure 15), we present the final execution plan to illustrate the structure of a complete plan. Due to the straightforward nature of the task, this figure does not include the reasoning process. For the superhero companions and fruit price ranking tasks (Figures 16 and 17), we emphasize the reasoning process and omit the block for the complete final plan for the sake of conciseness. The final plans for these tasks are similar in structure to the fruit placement task, essentially combining the substeps in the execution sequence at the bottom of the figures.
@@
B.4 Spatial Planning - Real Robots
B.4.1 Task Details
@@ B.4.2 Robot Setup
Appendix C Ablation Studies: Are All Parts of Wonderful Team Necessary?
@@
Figure 19 shows the workflow of the complete framework of Wonderful Team. We also provide the full prompt and example input and output corresponding to this workflow chart in Appendix C for more concrete details.
@@
Figure 19 shows the workflow of the complete framework of Wonderful Team. Detailed prompts, input examples, and output corresponding to this workflow can be found in Appendix C.
@@
C.1 Understanding What Each Part of Wonderful Team Does
@@
Appendix D Ablation Studies: VLLMs’ Spatial Reasoning Limitations and Potentials
D.1 Evaluating VLLM’s Spatial Understanding
@@ D.1.3 Deviation Analysis
D.2 Evaluating VLLMs’ Error Recognition and Correction
@@
D.2.1 Experimental Setup
To test this ability, we randomly generated 4 types of bounding boxes around the circle of interest. Examples are shown in 22. The types are:
@@ D.2.2 Results
From Table 7 and 8, we can see that GPT-4o demonstrated a very strong ability to examine and decide whether a bounding box is good enough just by visual inspection. This capability opens up new possibilities for self-refinements using current VLLMs. Even in cases where initial coordinate generation is not perfect, incorporating a checker as an additional layer of safety along the pipeline can iteratively improve coordinate accuracy until a satisfactory result is achieved.
@@
Appendix E Comparison with Other Methods
E.1 Replicating Natural Language as Policies Using gpt-4o
@@
E.2 Comparison with PIVOT
@@
  • PIVOT identifies a single action point for the target object, maintaining this as the goal throughout their iterative process. In contrast, our method offers two distinct workflows that the grounding manager can choose from before localization. When selecting an area point, such as a position between a box and a frame, we also employ point selection. However, for object selection, our method first identifies a center point, then determines a bounding box of appropriate size, and iteratively refines this bounding box until it is accurate. The grounding manager then selects an actionable point within the bounded area. We found that this bounding box process greatly enhances robustness and precision, especially for smaller objects or manipulation tasks that require more precise control. We further ablate and discuss this in Appendix E.2.
@@
In Table 10, we compare the experimental results of our method with those from PIVOT. While PIVOT originally utilizes GPT-4V in its framework, we implemented their approach using the more advanced GPT-4O to ensure a fair comparison. Our replication of their framework was carried out to the best of our knowledge to highlight the differences and performance improvements. Additionally, we include results obtained from their official HuggingFace demo to demonstrate the performance of their original implementation. For example output of different grounding approaches, please see 30.
@@
Parallel Runs: The original study also employs a parallel call strategy. To combine results from different runs, they explored two approaches: (1) fitting a new action distribution from the output actions and returning it, and (2) selecting a single best action using a VLM query. In our implementation, we used the second approach with “3 Iterations 3 Parallel“ combinations to enhance robustness in our comparison. Additionally, while the original implementation uses the same sampling radius for both width and height, we addressed this by defining separate radii for the shorter and longer edges of the input image. Grounding Team Only: Since PIVOT’s framework is primarily comparable to our grounding team, which focuses on processing object descriptions rather than broader tasks, we isolated the grounding component for a direct comparison with their method. Success Evaluation: For evaluation, we conducted 10 runs on different objects from a set of varied initial frames. A task was considered successful if the center point label of each target object had at least half of its area within the object’s boundary or if the center point fell within a specific range around the target area center, ensuring successful picking. Refer to caption Figure 29: Screenshot of HuggingFace PIVOT Demo Failure Mode Discussions It’s notable that PIVOT’s output on tabletop tasks does not over-perform the direct output from GPT-4o. However, this is with the help of the labeled coordinate system, which significantly enhances precision in quantification, as discussed in our motivation section. We further discuss the possible explanations of PIVOT failures: Incomplete Sampling Coverage: In 29, when attempting to select the left object, the initial sampling failed to provide sufficient coverage, with the majority of points being sampled from the center of the image and scattering on the purple paisley letter “V” instead of the target object with blue and purple stripes. As a result, subsequent iterations were confined to a suboptimal region, ultimately leading to poor final results. Difficulty in Recovery: During our implementation, we identified a critical limitation in the sampling strategy: if the sampling radius is too small, it becomes difficult to recover from an inadequate initial selection. Conversely, if the sampling radius is too large, the framework struggles to converge, as the sampled actions may scatter too broadly, reducing the effectiveness of the refinement process. Lack of Iterative Continuity: Another factor that may explain PIVOT’s low performance in precise location finding is the lack of continuity between iterations. Although the new set of actions is sampled from a distribution fitted using previously selected promising actions, there is a notable discontinuity in the process. For instance, if a good point is identified during one iteration, it is not guaranteed to be preserved in subsequent iterations. The framework’s fixed number of resampling processes means it cannot exit the process once a good point is found, potentially resulting in the loss of successful actions. This resampling process can lead to promising actions being either diluted or completely discarded in the next round due to inherent randomness, causing inefficiencies and inconsistencies as the framework may fail to build on previous successes. Messy Annotations: Additionally, the framework’s annotations can become cluttered, leading to a loss of crucial information from the original image. Unlike our approach, which maintains a clear connection to the original image to preserve full context, PIVOT’s method can lose track of the overall scene, making it difficult to refine action points effectively. This loss of context can be particularly detrimental in scenarios where precision and consistency are critical. Refer to caption Figure 30: Example Outputs - Wonderful Team vs PIVOT Point Selection vs. Bounding Box: Since the PIVOT method is inherently more similar to our area/point approach discussed earlier—where points are selected throughout the process without the aid of bounding boxes—we further compare PIVOT’s outputs with both our bounding box approach and our point approach. Figure 30 provides insight into how these methods perform relative to each other. While both PIVOT and our area/point approach can get reasonably close to the desired objects, they often lack the precision required for tasks involving small objects or when execution demands more accuracy than just proximity to the object. In Figure 31, we present example executions using the results from these methods. The task involves stacking the purple and blue striped letter “V” on top of the blue letter “V,“ followed by stacking the purple paisley letter “V” on top. For this execution, we used the PIVOT results from our implementation using gpt-4o, as the HuggingFace outputs were less reliable, with all points concentrated on the same object. The execution screenshots reveal that points not accurately placed on the object lead to failures in picking it up. On the bottom row of Figure 31, even though both points for the first pick-and-place action are technically correct, the misalignment causes the stacking task to partially fail, as the letters “V“ are not properly aligned, resulting in an unsuccessful stack. These results highlight the importance of considering whether a bounding box is needed in the iterative process. With the current level of visual reasoning skills in models, we found that incorporating a bounding box significantly enhances precision, reduces hallucinations, and adds robustness to the execution. Refer to caption Figure 31: Example Executions - Wonderful Team vs PIVOT
@@
E.3 Comparison with Language Models as Zero-Shot Trajectory Generators
E.3.1 Key Differences

### v5

@@ 6 Further Discussions
6.1 Comparison with Training-Based Methods
In recent years, the machine learning community has often seen new LLMs exceed the performance of previous generation fine-tuned models in zero-shot settings, despite the latter’s advantage of task-specific tuning. To explore this trend in the context of visual LLMs and robotics, we compare Wonderful Team with several methods that were at least partially fine-tuned on high-level planning robotics tasks.
@@ 6.2 Limitations: Where does Wonderful Team struggle?
High-Level Task Planning: We want to stress that we have only considered high level physical task planning in this paper. Robotics is an exceptionally broad area, and we can not make claims about the effectiveness of our methods on problems such as visual navigation or low-level kinematic control. We have also considered 3D manipulation environments exclusively, and do not make any claims about the applicability of these methods to other domains.
@@ Acknowledgments
  • Appendix B: Prompt and Implementation Details B
  • Appendix C: Experimental Details C
  • Appendix D: Ablations: Are All Parts of Wonderful Team Necessary? D
    • Understanding What Each Part of Wonderful Team Does D.1
  • Appendix E: Ablation Studies: VLLMs’ Spatial Reasoning Limitations and Potentials E
    • Evaluating VLLM’s Spatial Understanding E.1
    • Evaluating VLLMs’ Error Recognition and Correction E.2
  • Appendix F: Comparison with Other Methods F
    • Natural Language as Policies F.1
    • PIVOT F.2
    • Language Models as Zero-Shot Trajectory Generators F.3
    • MOKA F.4
@@ Appendix A Things are Moving Extremely Fast
In the Appendix E, we examine the impact of this rapid progress on the grounding team in particular, and show that older VLLMs often struggled to draw bounding boxes with any regularity, suggesting they lacked the fidelity needed for fine-grained robotic control.
Appendix B Prompt and Implementation Details
In this section, we detail the implementation of the Wonderful Team framework, focusing on the prompts used by the agents and their workflows. These prompts demonstrate how each agent fulfill specific roles, enabling the execution of complex tasks in unstructured environments. For reproducibility and further experimentation, the full prompts are available in our codebase. The Wonderful Team framework introduces a novel multi-agent VLLM-based approach for zero-shot high-level robotic planning, where specialized agents collaborate to decompose complex tasks into manageable components while ensuring cohesive objectives. This section outlines the overall architecture, agent roles, planning pipeline, and coordination mechanisms to provide a comprehensive understanding of the framework.
B.1 Multi-Agent Architecture
We adopt a multi-agent approach to address long-horizon, multi-step robotics tasks because traditional single-agent frameworks often suffer from scope overload, where a single agent must manage diverse task stages with different focus, leading to reduced effectiveness. By dividing the problem into smaller, specialized sub-tasks, each agent focuses on a specific role, such as planning, grounding, or intermediate validation, ensuring more precise and efficient handling of each step. Each agent also stores their previous input and output to their own internal memory to maintain context-awareness and scope consistency when working on tasks. The agents collaborate by sharing information and verifying intermediate outputs, which minimizes errors and enhances robustness. With this multi-agent approach, we highlight several characteristics:
  • Specialized functionality: Dividing responsibilities among agents tailored to specific roles ensures efficient task execution, as no single agent is burdened with managing every aspect of the process.
  • Internal memory: Agents equipped with memory capabilities can make consistent, context-aware decisions, leading to higher-quality outcomes.
  • Different scope of information: Assigning different scopes to each agent allows them to concentrate on their specific objectives without being distracted by irrelevant details, such as a high-level agent focusing on strategic planning without managing low-level execution and vice versa.
B.2 Wonderful Team Pipeline
Our framework operates through three primary stages: high-level planning, coordinate-level target location grounding, and low-level action generation. A memory agent collects outputs of other agents across all stages and maintains a system memory of important task-relevant information.
B.2.1 High-Level Planning
The high-level planning phase involves iterative refinement between the Supervisor and Verification agents, as outlined in Algorithm 1.
Input: Task Prompt (text, image, or both), Environment Observation (image)
Output: Initial Plan, Target Subgoal Object/Area
The Supervisor focuses on high-level objectives and proposes plans with subgoals. The Verification agent evaluates the plan for feasibility and logic, checking for issues such as collision risks, physical constraints, and missing prerequisites. In each iteration, the Verification agent raises specific concerns about subgoals and tracks its feedback and completed checks using internal memory to avoid redundancy. The Supervisor revises the plan based on the Verification agent’s feedback. This process repeats until the Verification agent approves the plan, ensuring it meets all constraints.
Table 2: A simplified prompt for Supervisor agent proposing high-level plans.
You are the supervisor creating and modifying a robotics plan.
Your focus should be on overall task-level objectives, providing guidelines and direction for other agents. Low-level grounding and control details are not within your scope of concern.
Here is the task: {task prompt}
Here is the environment: {env}
(If feedback is applicable) Here are the questions raised by the verification agent on potential issues based on your previous plan.
Please output a high-level plan consisting of a list of subgoals.
Table 3: A simplified prompt for Verification agent verifying high-level plans.
You are a verification agent responsible for checking the feasibility of a robotics task plan and giving feedback for revision if applicable.
Your sole objective is to ensure the plan itself is logical and achievable.
Here is the current plan: {plan}
Here is the environment: {env}
Check for these potential issues:
1. Collision avoidance
2. Physical constraints
3. Missing prerequisite steps
Please output APPROVED or feedback for plan revision.
Algorithm 1 High-Level Planning
1:Task prompt $T$, Environment observation $E$
2:Verified plan $P$, Target objects/areas list $L$
3:$P\leftarrow\text{Supervisor.CreatePlan}(T,E)$
4:while not approved do
[…]
6:     if feedback is approved then
7:         break
8:     else
9:         $P\leftarrow\text{Supervisor.Revise}(P\mid\text{feedback},E)$
10:     end if
11:end while
12:$L\leftarrow\text{Supervisor.ExtractTargets}(P)$
13:return $P,L$
B.2.2 Coordinate-Level Target Location Grounding
The Grounding Team has three agents: the Manager, the Mover, and the Checker, to identify action points for targets in an environment based on a verified plan. For each target, the Manager observe the entire environment image and initializes a bounding box. To streamline their tasks, the Mover and Checker agents are provided with a zoomed-in image centered on the bounding box rather than the entire environment image, minimizing distractions from irrelevant details. The Checker then evaluates the bounding box image for feedback and approval. If the bounding box is not approved, the Mover revises the bounding box based on feedback, and the process repeats until approval is obtained. Once approved, the Manager determines an action point within the bounding box, and it is added to the final set of action points. Details are shown in Algorithm 2.
Table 4: A simplified prompt for Grounding Team Manager agent estimating initial bounding boxes.
You are the manager of the grounding team.
With full visibility of the scope, tasks, and subgoals, your task is to guide two workers who refine your approximate guesses of target locations into precise bounding boxes using zoomed-in views. When the two workers have finalized the box, you choose an action point based on the task and plan, considering context like pushing in a direction or picking up an object.
Here is the environment: {env}
Here is the verified plan: {plan}
Here is the current object/area we are working on: {target}
Please give an approximate starting point and bounding box.
Table 5: A simplified prompt for Grounding Team Mover agent adjusting bounding boxes.
You are an agent responsible for adjusting bounding box location and dimensions, ensuring bounding boxes are accurately positioned and sized for precise object manipulation tasks.
Your adjustments must be meaningful and significant before the final version is concluded.
Here is the object you are working on: {object}
Here is the current bounding box: {bounding box specification}
Here is the object: {zoomed-in image around current bounding box}
Please propose a revision.
Table 6: A simplified prompt for Grounding Team Checker agent evaluating bounding boxes.
You are responsible for verifying the accuracy and alignment of bounding boxes for objects of interest.
Collaborating with another agent who proposes revisions, you evaluate whether the suggested changes improve the current bounding box. Additionally, you decide when no more adjustments are needed for a bounding box for final output
Here is the object you are working on: {object}
Here is the revision: {bounding box before and after revision}
Please determine if the revision for the bounding box is acceptable; if it is acceptable, you have to then decide if it no longer needs further revision.
Table 7: A simplified prompt for Grounding Team Manager agent selecting action points.
Your task is to analyze the image and task requirements to identify an appropriate point of action based on the task’s characteristics. Ensure the chosen point aligns with the plan objectives and key object properties.
Here is the object: {zoomed-in environment}
Here is the revised plan: {plan}
Here is the current object/area we are working on: {target}
Please give the action point.
Algorithm 2 Coordinate-Level Target Location Grounding
1:Target objects/areas $L$, Environment $E$, Verified Plan $P$
2:Action points $A$
3:for target $t$ in $L$ do
4:     $bbox_{t}\leftarrow\text{Manager.InitializeBox}(t,E,P)$
5:     while not approved do
[…]
7:         if feedback is approved then
8:              break
9:         end if
10:         $bbox_{t}\leftarrow\text{Mover.Revise}(bbox_{t})$
11:     end while
12:     $a_{t}\leftarrow\text{Manager.DetermineActionPoint}(bbox_{t},P)$
13:     $A\leftarrow A\cup\{a_{t}\}$
14:end for
15:return $A$
B.2.3 Low-Level Action Generation
In the final stage, the Supervisor agent consults the system memory maintained by the memory agent and converts the verified plan $P$ and action points $A$ into low-level commands. These commands form the executable action sequence, ensuring alignment with the spatial and temporal constraints in the plan and grounding information. The Supervisor is tasked with this step because each agent operates within its own scope and maintains an internal memory of its specific inputs and outputs. Other agents are focused on their specialized roles and responsibilities unrelated to the overall task objectives. In contrast, the Supervisor has been consistently managing the high-level task throughout the process, ensuring continuity and focus. With the finalized grounding information readily available, the Supervisor can efficiently generate the action sequence based on the plan it previously produced, ensuring no critical constraints or details from the system memory are lost.
Table 8: A simplified prompt for Supervisor agent generating actions.
You have access to a system memory containing all essential task-relevant information summarized from the multi-agent workflow, along with a top-view image of the environment. Your task is to translate this information into a sequence of actions that align with the plan subgoals.
Here is the memory: {memory}
Here is the environment: {env}
Please convert to the final plan.
Appendix C Experimental Details
C.1 Evaluation Protocol
@@
C.2 Multimodal Reasoning - Simulated VIMABench
@@ C.2.1 Task Details
C.2.2 Full Experimental Results
@@
Task Num VIMA 200M Instruct2Act NLaP (w/o CoT) NLaP TG MOKA Ours
1: Visual Manipulation 99 91 93 100 60 20 100
2: Scene Understanding 100 81 60 67 40 20 100
3: Rotate 100 98 93 93 80 0 100
4: Rearrange 97 79 52 73 - 0 80
5: Rearrange then Restore 54.5 72 25 73 - 0 70
6: Novel Adjective 100 82 13 43 10 10 70
7: Novel Noun 99 88 8 80 0 10 100
*8: Novel Adjective and Noun - - - - - 20 60
*9: Twist 17.5 - - - - 0 50
10: Follow Motion - 35 0 12 - 0 10
11: Follow Order 90.5 72 0 0 - 0 0
12: Without Exceeding 93 68 17 47 10 0 90
*13: Without Touching - 0 0 3 - 0 40
14: Same Texture - 80 3 71 - 0 100
15: Same Shape 97.5 78 10 80 0 0 100
16: Manipulate Old Neighbor 46 64 8 20 0 0 90
17: Pick in Order then Restore 43.5 85 10 30 0 0 90
Note: Tasks marked with a star (*) were excluded from Figure 10(c), but their results are included in this table for completeness.
Reasons for Task Exclusion in Figure 10(c): Tasks marked with a star (*) were excluded from the main paper’s results for the following reasons:
1. Nature of Tasks: Certain categories, such as Visual Goal Reaching (Tasks 4 and 5) and One-shot Video Imitation (Tasks 10 and 11), were excluded because they are not ideal for evaluating Vision-Language Learning Model (VLLM) capabilities without additional task-specific prompting.
For example, Figure 14 illustrates Task 11 in the One-shot Video Imitation category, where several consecutive frames serve as “goal scenes.” Without task-specific prompting or training, it becomes challenging to infer the required actions between frames, as there is no single definitive solution.
@@
Consider the transition from Frame 1 to Frame 2 in the example above: - One possible approach involves moving the yellow “O” onto the red “O.” - Alternatively, another approach might first remove the red “O” and then place the yellow “O” in the same position.
For methods capable of processing multimodal prompts, these tasks require additional tools or workflows, such as detailed explanations of the relationships between consecutive frames (e.g., treating them as a continuous video with specific temporal assumptions). This adds complexity to zero-shot evaluation, making task-specific prompting necessary to infer inter-frame relationships. While task-specific prompting could improve performance, it falls outside the scope of our research. Consequently, evaluations presented in Table 9 were conducted using standardized prompts.
For methods that do not support multimodal inputs, textual prompts for these tasks are often insufficient. They lack the spatial and temporal context necessary to infer relationships, making it difficult to interpret frame-based tasks correctly.
Tasks requiring such inferences include Tasks 4, 5, 9, 10, 11, and 17.
2. Missing Baseline Results: Tasks 8, 9, 13, and 14 were excluded due to the absence of baseline results for comparison.
A comprehensive list of tasks, along with video illustrations, is available at this link.
C.3 Implicit Goal Inference - Real Robots
C.3.1 Task Details
@@ C.3.2 Robot Setup
C.3.3 Results
Our real robot experiments demonstrated that our framework successfully completed all three tasks 100% of the time. Note that we did not modify any of the prompt or pipeline moving from simulated VIMABench environment to the tasks on the real robot. It was surprising to us how robust the reasoning and planning capabilities of Wonderful Team are. This section provides qualitative results from these experiments, illustrated in Figures 16, 17, and 18. These figures highlight specific aspects of the tasks, illustrating the effectiveness of our framework. It is important to note that these results only reflect the work of the planning team. The role of the grounding team, locating objects and determining their positions, is crucial for the successful execution of these plans.
@@ C.3.3 Results
In the fruit placement task (Figure 16), we present the final execution plan to illustrate the structure of a complete plan. Due to the straightforward nature of the task, this figure does not include the reasoning process. For the superhero companions and fruit price ranking tasks (Figures 17 and 18), we emphasize the reasoning process and omit the block for the complete final plan for the sake of conciseness. The final plans for these tasks are similar in structure to the fruit placement task, essentially combining the substeps in the execution sequence at the bottom of the figures.
@@
C.4 Spatial Planning - Real Robots
C.4.1 Task Details
@@ C.4.2 Robot Setup
Appendix D Ablation Studies: Are All Parts of Wonderful Team Necessary?
@@
Figure 20 shows the workflow of the complete framework of Wonderful Team. We also provide the full prompt and example input and output corresponding to this workflow chart in Appendix D for more concrete details.
@@
Figure 20 shows the workflow of the complete framework of Wonderful Team. Detailed prompts, input examples, and output corresponding to this workflow can be found in Appendix D.
@@
D.1 Understanding What Each Part of Wonderful Team Does
@@
Appendix E Ablation Studies: VLLMs’ Spatial Reasoning Limitations and Potentials
E.1 Evaluating VLLM’s Spatial Understanding
@@ E.1.3 Deviation Analysis
E.2 Evaluating VLLMs’ Error Recognition and Correction
@@
E.2.1 Experimental Setup
To test this ability, we randomly generated 4 types of bounding boxes around the circle of interest. Examples are shown in 23. The types are:
@@ E.2.2 Results
From Table 14 and 15, we can see that GPT-4o demonstrated a very strong ability to examine and decide whether a bounding box is good enough just by visual inspection. This capability opens up new possibilities for self-refinements using current VLLMs. Even in cases where initial coordinate generation is not perfect, incorporating a checker as an additional layer of safety along the pipeline can iteratively improve coordinate accuracy until a satisfactory result is achieved.
@@
Appendix F Comparison with Other Methods
F.1 Replicating Natural Language as Policies Using gpt-4o
@@
F.2 Comparison with PIVOT
@@
  • PIVOT identifies a single action point for the target object, maintaining this as the goal throughout their iterative process. In contrast, our method offers two distinct workflows that the grounding manager can choose from before localization. When selecting an area point, such as a position between a box and a frame, we also employ point selection. However, for object selection, our method first identifies a center point, then determines a bounding box of appropriate size, and iteratively refines this bounding box until it is accurate. The grounding manager then selects an actionable point within the bounded area. We found that this bounding box process greatly enhances robustness and precision, especially for smaller objects or manipulation tasks that require more precise control. We further ablate and discuss this in Appendix F.2.
@@
In Table 17, we compare the experimental results of our method with those from PIVOT. While PIVOT originally utilizes GPT-4V in its framework, we implemented their approach using the more advanced GPT-4O to ensure a fair comparison. Our replication of their framework was carried out to the best of our knowledge to highlight the differences and performance improvements. Additionally, we include results obtained from their official HuggingFace demo to demonstrate the performance of their original implementation. For example output of different grounding approaches, please see 31.
@@
Parallel Runs: The original study also employs a parallel call strategy. To combine results from different runs, they explored two approaches: (1) fitting a new action distribution from the output actions and returning it, and (2) selecting a single best action using a VLM query. In our implementation, we used the second approach with “3 Iterations 3 Parallel“ combinations to enhance robustness in our comparison. Additionally, while the original implementation uses the same sampling radius for both width and height, we addressed this by defining separate radii for the shorter and longer edges of the input image. Grounding Team Only: Since PIVOT’s framework is primarily comparable to our grounding team, which focuses on processing object descriptions rather than broader tasks, we isolated the grounding component for a direct comparison with their method. Success Evaluation: For evaluation, we conducted 10 runs on different objects from a set of varied initial frames. A task was considered successful if the center point label of each target object had at least half of its area within the object’s boundary or if the center point fell within a specific range around the target area center, ensuring successful picking. Refer to caption Figure 30: Screenshot of HuggingFace PIVOT Demo Failure Mode Discussions It’s notable that PIVOT’s output on tabletop tasks does not over-perform the direct output from GPT-4o. However, this is with the help of the labeled coordinate system, which significantly enhances precision in quantification, as discussed in our motivation section. We further discuss the possible explanations of PIVOT failures: Incomplete Sampling Coverage: In 30, when attempting to select the left object, the initial sampling failed to provide sufficient coverage, with the majority of points being sampled from the center of the image and scattering on the purple paisley letter “V” instead of the target object with blue and purple stripes. As a result, subsequent iterations were confined to a suboptimal region, ultimately leading to poor final results. Difficulty in Recovery: During our implementation, we identified a critical limitation in the sampling strategy: if the sampling radius is too small, it becomes difficult to recover from an inadequate initial selection. Conversely, if the sampling radius is too large, the framework struggles to converge, as the sampled actions may scatter too broadly, reducing the effectiveness of the refinement process. Lack of Iterative Continuity: Another factor that may explain PIVOT’s low performance in precise location finding is the lack of continuity between iterations. Although the new set of actions is sampled from a distribution fitted using previously selected promising actions, there is a notable discontinuity in the process. For instance, if a good point is identified during one iteration, it is not guaranteed to be preserved in subsequent iterations. The framework’s fixed number of resampling processes means it cannot exit the process once a good point is found, potentially resulting in the loss of successful actions. This resampling process can lead to promising actions being either diluted or completely discarded in the next round due to inherent randomness, causing inefficiencies and inconsistencies as the framework may fail to build on previous successes. Messy Annotations: Additionally, the framework’s annotations can become cluttered, leading to a loss of crucial information from the original image. Unlike our approach, which maintains a clear connection to the original image to preserve full context, PIVOT’s method can lose track of the overall scene, making it difficult to refine action points effectively. This loss of context can be particularly detrimental in scenarios where precision and consistency are critical. Refer to caption Figure 31: Example Outputs - Wonderful Team vs PIVOT Point Selection vs. Bounding Box: Since the PIVOT method is inherently more similar to our area/point approach discussed earlier—where points are selected throughout the process without the aid of bounding boxes—we further compare PIVOT’s outputs with both our bounding box approach and our point approach. Figure 31 provides insight into how these methods perform relative to each other. While both PIVOT and our area/point approach can get reasonably close to the desired objects, they often lack the precision required for tasks involving small objects or when execution demands more accuracy than just proximity to the object. In Figure 32, we present example executions using the results from these methods. The task involves stacking the purple and blue striped letter “V” on top of the blue letter “V,“ followed by stacking the purple paisley letter “V” on top. For this execution, we used the PIVOT results from our implementation using gpt-4o, as the HuggingFace outputs were less reliable, with all points concentrated on the same object. The execution screenshots reveal that points not accurately placed on the object lead to failures in picking it up. On the bottom row of Figure 32, even though both points for the first pick-and-place action are technically correct, the misalignment causes the stacking task to partially fail, as the letters “V“ are not properly aligned, resulting in an unsuccessful stack. These results highlight the importance of considering whether a bounding box is needed in the iterative process. With the current level of visual reasoning skills in models, we found that incorporating a bounding box significantly enhances precision, reduces hallucinations, and adds robustness to the execution. Refer to caption Figure 32: Example Executions - Wonderful Team vs PIVOT
@@
F.3 Comparison with Language Models as Zero-Shot Trajectory Generators
F.3.1 Key Differences